# Personalized Spiking Neural Networks with Ferroelectric Synapses for EEG Signal Processing

### v1

@@ Abstract
Electroencephalography (EEG)-based brain–computer interfaces (BCIs) are strongly affected by non-stationary neural signals that vary across sessions and individuals, limiting the generalization of subject-agnostic models and motivating adaptive and personalized learning on resource-constrained platforms. Programmable memristive hardware offers a promising substrate for such post-deployment adaptation; however, practical realization is challenged by limited weight resolution, device variability, nonlinear programming dynamics, and finite device endurance. In this work, we show that Spiking Neural Networks (SNNs) can be deployed on ferroelectric memristive synaptic devices for adaptive EEG-based motor imagery decoding under realistic device constraints. We fabricate, characterize, and model the weight update rule in ferroelectric synapses. We evaluate a convolutional–recurrent SNN architecture under two complementary deployment strategies: (i) device-aware training using a ferroelectric synapse model, and (ii) transfer of software-trained weights followed by low-overhead on-device re-tuning. To enable efficient adaptation, we introduce a device-aware weight-update strategy in which gradual gradient-based updates are accumulated digitally and converted into discrete programming events only when a threshold is exceeded, emulating the nonlinear, state-dependent programming dynamics of the ferroelectric device while significantly reducing programming frequency. Both deployment strategies achieve classification performance comparable to state-of-the-art software-based SNNs. Furthermore, subject-specific transfer learning achieved by retraining only the final network layers leads to improved classification accuracy. These results demonstrate that programmable ferroelectric hardware can support robust, low-overhead adaptation in spiking neural networks, opening a practical path toward personalized neuromorphic processing of non-stationary neural signals.
@@ 1 Introduction
Brain–computer interfaces (BCIs) based on electroencephalography (EEG) have attracted significant interest due to their potential for assistive communication in patients with severe motor impairments [bamdad2015application, kawala2021summary, kanungo2021wheelchair]. From a deployment perspective, practical EEG decoding must operate under stringent constraints on latency and power: decisions should be produced with minimal delay after the user’s intent while maintaining ultra-low energy consumption to ensure acceptable battery lifetime and to limit tissue heating in wearable and implantable settings. More broadly, wearable sensors and implantable devices are increasingly pushing artificial intelligence workloads toward the extreme edge [wang2017flexible, shi2016edge, rancea2024edge], where strict thermal and energy budgets fundamentally shape algorithm and hardware choices [vazquez2021compute]. In this context, EEG-based decoding represents a representative extreme-edge signal-processing workload, requiring both low end-to-end latency and high energy efficiency.
@@
Within this architectural framework, beyond-CMOS memories have emerged as promising building blocks for in-memory computing [theis2017end, mehonic2024roadmap]. In resistive memory arrays, physical laws such as Ohm’s and Kirchhoff’s laws can be exploited to perform multiply-and-accumulate operations directly within memory, reducing data transfer and its associated energy overhead. Memristive devices can emulate synaptic functionality by encoding network weights as programmable conductance states connecting successive layers. Their nonvolatile retention and sub-femtojoule read energy [covi2022challenges], together with demonstrated multi-level programmability [thomas2024versatile], make them attractive for adaptive processing of non-stationary physiological signals [fang2022compact]. Analog or multi-state behavior suitable for learning has been demonstrated across material systems, including phase-change and valence-change memories [rao2023thousands]. Ferroelectric memristive synapses, which consist of an ultra-thin ferroelectric film (zirconium doped hafnium oxide) between two asymmetric electrodes, further exhibit high endurance (above 110 cycles [begon2022scaled]) and co-integration to CMOS [begon-lours_back-end--line_2024].
However, translating these motivations into practical EEG decoding systems remains challenging. Memristive synapses are constrained by limited effective weight resolution, device-to-device variability, nonlinear and state-dependent programming dynamics, and finite endurance, all of which can degrade accuracy and stability when deploying learned models on hardware. At the system level, two common deployment strategies are (i) offline learning in software followed by weight transfer through device programming, and (ii) direct on-device learning using local update rules [alibart2013pattern]. Both strategies must contend with the highly non-stationary nature of EEG and the significant variability across trials, recording sessions, and individuals, which limits the generalization of subject-agnostic models and motivates adaptive and personalized learning. At the same time, personalization [zheng2016personalizing] must be achieved under strict constraints on energy consumption, programming frequency, and device lifetime, requiring learning mechanisms that are robust to device non-idealities while enabling low-overhead post-deployment adaptation.
@@ 2.1 Motor imagery dataset
In this study, we used the Physionet [goldberger2000physiobank] motor imagery dataset [schalk2004bci2000]. This database comprises recording from 109 participants through a 64 channel EEG headset using BCI2000 system with a sampling rate of 160 Hz. These recordings were obtained from participants engaged in a sequence of motor and imagery activities. The tasks related to action are not considered in this study, because the model should be trained on people not able to perform actions. Same as the study by [kumar2022decoding], data from six subjects were removed, resulting in 103 subjects used in our study. The recordings are classified in two subsets of imagined actions and left hand imagery task contained 2336 trials in total and right hand imagery task contained 2299 trials in total. The signals were band-pass filtered between 0.1-80 Hz, and a 1-second segment (0–1 s after cue) was extracted for each trial.
@@ 2.4 Synaptic device
A ferroelectric synaptic device with the same functional stack from that of of ref. [begon-lours_back-end--line_2024] was fabricated on a silicon subtrate. The substrate is a 0.8 µm thick thermal SiO2 on Si chip. Using Atomic Layer Deposition (ALD), 20 nm of TiN was deposited at 300°C, then 45 cycles of WOx at 360°C. Afterwards, a [HfO2/ZrO2] nanolaminate is deposited at 300°C consisting of five supercycles, each comprising five cycles with tetrakis (ethylmethylamino) hafnium (IV) and O2, and ten cycles with bis (methylcyclopentadienyl) (methyl) (methoxy) zirconium (IV) and O2. Ten nanometers of TiN were deposited in situ. Crystallization was performed with a millisecond flash lamp anneal: the sample was preheated at 450°C, followed by a 20 ms pulse of 90 J/cm2. A 50 nm thick W metal electrode was then sputtered. The top electrode was defined by UV lithography and reactive ion etching (RIE). A 100 nm thick SiO2 passivation layer was then sputtered. Vias to the device’s top electrode contacts were defined by UV lithography. The SiO2 layer was etched by RIE. A cross-section is represented in Fig. 2(a).
@@
To characterize multiple conductance states, a 8300 µm2 device was repeatedly cycled between high and low-resistance states (HRS/LRS) applying a sequence of programming pulses, as shown in Fig. 2(b). The programming pulses had a constant width of 50 µs and their amplitude gradually increased from 0 V to +1 V for potentiation and from 0 V to –1.25 V for depression, in +/-50 mV steps. For read operations, bipolar triangular pulses of +/-100 mV were applied, and the resulting current was measured using a source–measure unit (SMU). The measured resistance values ranged from approximately 10 MΩ to 100 MΩ. The physical mechanisms leading to a change in conductance upon programming pulses are detailed in ref.[begon-lours_back-end--line_2024]: at each pulse, the direction of the polarization for a fraction of the ferroelectric domains flips. The screening of the polarization results in a local redistribution of charges, modifying the conduction through the insulator.
The programmable conductance states were grouped according to the magnitude of the programming pulses, as shown in Fig. 2(c–d) for long-term depression (LTD) and long-term potentiation (LTP).
@@ 2.7 On-device learning
During online learning, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ using backpropagated gradients (Fig. 4). The hardware update is not performed as long as $|\delta w_{\mathrm{acc}}|<\varepsilon$, where $\varepsilon$ is a threshold that represents the minimum change that can be reliably programmed in the device.
@@ 4 Conclusion
This work demonstrates that programmable memristive hardware can support adaptive spiking neural networks for EEG-based brain–computer interfaces, enabling learning directly within a memristive synapse model. Network accuracy is largely preserved when transitioning from an ideal digital implementation to a memristive realization with nonlinear and asymmetric update dynamics, indicating that the proposed learning framework is tolerant to key device-level non-idealities. These results support the practical feasibility of deploying spiking neural networks on memristive neuromorphic hardware. Our results further highlight that such SNNs are robust to limited weight resolution and programming noise, as evidenced by the quantization and additive noise experiments. Importantly, we demonstrate that residual performance degradation can be recovered with only a few on-device learning epochs, mitigating the need for extensive reprogramming and thereby contributing to improved device lifetime.
@@
The PhysioNet EEG dataset used in this study is publicly available at https://physionet.org/content/eegmmidb/1.0.0/. The data presented in this paper are available from the authors upon reasonable request.

### v2

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) are strongly affected by non-stationary neural signals that vary across sessions and individuals, limiting the generalization of subject-agnostic models and motivating adaptive and personalized learning on resource-constrained platforms. Programmable memristive hardware offers a promising substrate for such post-deployment adaptation; however, practical realization is challenged by limited weight resolution, device variability, nonlinear programming dynamics, and finite device endurance. In this work, we show that spiking neural networks (SNNs) can be deployed on ferroelectric memristive synaptic devices for adaptive EEG-based motor imagery decoding under realistic device constraints. We fabricate, characterize, and model ferroelectric synapses. We evaluate a convolutional-recurrent SNN architecture under two complementary deployment strategies: (i) device-aware training using a ferroelectric synapse model, and (ii) transfer of software-trained weights followed by low-overhead on-device re-tuning. To enable efficient adaptation, we introduce a device-aware weight-update strategy in which gradient-based updates are accumulated digitally and converted into discrete programming events only when a threshold is exceeded, emulating nonlinear, state-dependent programming dynamics while reducing programming frequency. Both deployment strategies achieve classification performance comparable to state-of-the-art software-based SNNs. Furthermore, subject-specific transfer learning achieved by retraining only the final network layers improves classification accuracy. These results demonstrate that programmable ferroelectric hardware can support robust, low-overhead adaptation in spiking neural networks, opening a practical path toward personalized neuromorphic processing of neural signals.
@@ 1 Introduction
Brain-computer interfaces (BCIs) based on electroencephalography (EEG) have attracted significant interest due to their potential for assistive communication in patients with severe motor impairments [bamdad2015application, kawala2021summary, kanungo2021wheelchair]. From a deployment perspective, practical EEG decoding must operate under stringent constraints on latency and power: decisions should be produced with minimal delay after the user’s intent while maintaining ultra-low energy consumption to ensure acceptable battery lifetime and to limit tissue heating in wearable and implantable settings. More broadly, wearable sensors and implantable devices are increasingly pushing artificial intelligence workloads toward the extreme edge [wang2017flexible, shi2016edge, rancea2024edge], where strict thermal and energy budgets fundamentally shape algorithm and hardware choices [vazquez2021compute]. In this context, EEG-based decoding represents a representative extreme-edge signal-processing workload, requiring both low end-to-end latency and high energy efficiency.
@@
Within this architectural framework, beyond-CMOS memories have emerged as promising building blocks for in-memory computing [theis2017end, mehonic2024roadmap]. In resistive memory arrays, physical laws such as Ohm’s and Kirchhoff’s laws can be exploited to perform multiply-and-accumulate operations directly within memory, reducing data transfer and its associated energy overhead. Memristive devices can emulate synaptic functionality by encoding network weights as programmable conductance states connecting successive layers. Their nonvolatile retention and sub-femtojoule read energy [covi2022challenges], together with demonstrated multi-level programmability [thomas2024versatile], make them attractive for adaptive processing of non-stationary physiological signals [fang2022compact]. Analog or multi-state behavior suitable for learning has been demonstrated across material systems, including phase-change and valence-change memories [rao2023thousands]. Ferroelectric memristive synapses, which consist of an ultra-thin ferroelectric film (zirconium doped hafnium oxide) between two asymmetric electrodes, further exhibit high endurance (above 110 cycles [begon2022scaled]) and co-integration to CMOS [begon2024back].
However, translating these motivations into practical EEG decoding systems remains challenging. Memristive synapses are constrained by limited effective weight resolution, device-to-device variability, nonlinear and state-dependent programming dynamics, and finite endurance, all of which can degrade accuracy and stability when deploying learned models on hardware. At the system level, two common deployment strategies are (i) offline learning in software followed by weight transfer through device programming, and (ii) direct on-device learning using local update rules [garg2025unsupervised, alibart2013pattern]. Both strategies must contend with the highly non-stationary nature of EEG and the significant variability across trials, recording sessions, and individuals, which limits the generalization of subject-agnostic models and motivates adaptive and personalized learning. At the same time, personalization [zheng2016personalizing] must be achieved under strict constraints on energy consumption, programming frequency, and device lifetime, requiring learning mechanisms that are robust to device non-idealities while enabling low-overhead post-deployment adaptation.
@@ 2.1 Motor imagery dataset
In this study, we used the Physionet [goldberger2000physiobank] motor imagery dataset [schalk2004bci2000]. This database comprises recording from 109 participants through a 64 channel EEG headset using BCI2000 system with a sampling rate of 160 Hz. These recordings were obtained from participants engaged in a sequence of motor and imagery activities. The tasks related to action are not considered in this study, because the model should be trained on people not able to perform actions. Same as the study by [kumar2022decoding], data from six subjects were removed, resulting in 103 subjects used in our study. The recordings are classified in two subsets of imagined actions and left hand imagery task contained 2336 trials in total and right hand imagery task contained 2299 trials in total. The signals were band-pass filtered between 0.1-80 Hz, and a 1-second segment (0-1 s after cue) was extracted for each trial.
@@ 2.4 Synaptic device
A ferroelectric synaptic device with the same functional stack as that of [begon2024back] was fabricated on a silicon subtrate. The substrate is a 0.8 µm thick thermal SiO2 on Si chip. Using Atomic Layer Deposition (ALD), 20 nm of TiN was deposited at 300°C, then 45 cycles of WOx at 360°C. Afterwards, a [HfO2/ZrO2] nanolaminate is deposited at 300°C consisting of five supercycles, each comprising five cycles with tetrakis (ethylmethylamino) hafnium (IV) and O2, and ten cycles with bis (methylcyclopentadienyl) (methyl) (methoxy) zirconium (IV) and O2. Ten nanometers of TiN were deposited in situ. Crystallization was performed with a millisecond flash lamp anneal: the sample was preheated at 450°C, followed by a 20 ms pulse of 90 J/cm2. A 50 nm thick W metal electrode was then sputtered. The top electrode was defined by UV lithography and reactive ion etching (RIE). A 100 nm thick SiO2 passivation layer was then sputtered. Vias to the device’s top electrode contacts were defined by UV lithography. The SiO2 layer was etched by RIE. A cross-section is represented in Fig. 2(a).
@@
To characterize multiple conductance states, a 8300 µm2 device was repeatedly cycled between high and low-resistance states (HRS/LRS) applying a sequence of programming pulses, as shown in Fig. 2(b). The programming pulses had a constant width of 50 µs and their amplitude gradually increased from 0 V to +1 V for potentiation and from 0 V to -1.25 V for depression, in +/-50 mV steps. For read operations, bipolar triangular pulses of +/-100 mV were applied, and the resulting current was measured using a source-measure unit (SMU). The measured resistance values ranged from approximately 10 MΩ to 100 MΩ. The physical mechanisms leading to a change in conductance upon programming pulses are detailed in ref.[begon2024back]: at each pulse, the direction of the polarization for a fraction of the ferroelectric domains flips. The screening of the polarization results in a local redistribution of charges, modifying the conduction through the insulator.
The programmable conductance states were grouped according to the magnitude of the programming pulses, as shown in Fig. 2(c-d) for long-term depression (LTD) and long-term potentiation (LTP).
@@ 2.7 On-device learning
During online learning, each synaptic weight maintains an accumulated update $\delta w_{\mathrm{acc}}$ using backpropagated gradients (Fig. 4). The hardware update is not performed as long as $|\delta w_{\mathrm{acc}}|<\varepsilon$, where $\varepsilon$ is a threshold that represents the minimum change that can be reliably programmed in the device. This accumulation-based update mechanism is conceptually similar to the mixed-precision learning strategy previously proposed for non-spiking networks with memristive [nandakumar2020mixed] and ferroelectric synapses [garg2025energy].
@@ 4 Conclusion
This work demonstrates that programmable memristive hardware can support adaptive spiking neural networks for EEG-based brain-computer interfaces, enabling learning directly within a memristive synapse model. Network accuracy is largely preserved when transitioning from an ideal digital implementation to a memristive realization with nonlinear and asymmetric update dynamics, indicating that the proposed learning framework is tolerant to key device-level non-idealities. These results support the practical feasibility of deploying spiking neural networks on memristive neuromorphic hardware. Our results further highlight that such SNNs are robust to limited weight resolution and programming noise, as evidenced by the quantization and additive noise experiments. Importantly, we demonstrate that residual performance degradation can be recovered with only a few on-device learning epochs, mitigating the need for extensive reprogramming and thereby contributing to improved device lifetime.
@@
The PhysioNet EEG dataset used in this study is publicly available at https://physionet.org/content/
eegmmidb/1.0.0/.